40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 7 10:26:36 2018, keerthana, Update, PSL, observing the resonance signal corresponding to the injected frequency. photodetector_alignment.jpgdesign1.PNG
    Reply  Thu Jun 7 22:36:09 2018, not keerthana, Update, PSL, observing the resonance signal corresponding to the injected frequency. 
    Reply  Fri Jun 8 00:36:54 2018, gautam, Update, PSL, observing the resonance signal corresponding to the injected frequency. 
    Reply  Fri Jun 8 01:08:22 2018, johannes, Update, PSL, First light of AUX at YEND AG4395A_07-06-2018_205019.pdfPSL_AUX_SETUP.pdfAS_AUX_SETUP.pdf
       Reply  Mon Jun 18 14:32:42 2018, gautam, Update, PSL, Optics on AS table 
          Reply  Mon Jun 18 15:59:17 2018, johannes, Update, PSL, Optics on AS table 
             Reply  Mon Jun 18 16:57:54 2018, Koji, Update, PSL, Optics on AS table 
             Reply  Fri Jun 22 14:18:04 2018, Steve, Update, PSL, Optics on AS table 
Message ID: 13924     Entry time: Thu Jun 7 10:26:36 2018     Reply to this: 13930   13931   13932
Author: keerthana 
Type: Update 
Category: PSL 
Subject: observing the resonance signal corresponding to the injected frequency. 

(Johannes, Koji, Keerthana)

The PLL loop ensures that the frequency difference between the PSL laser and the AUX laser is equal to the frequency we provide to the Local Oscillator (LO) with the help of a Marconi. Only a small pick off part of both the AUX and PSL lasers are going to the PLL loop. The other part of both the lasers are going to the interferometer. Before entering into the optical fibre, the AUX laser passes through an AOM which changes its frequency by an amount of 80MHz. When the PLL is locked, the frequency coming out of the PLL will be equal to the frequency set up in the Marconi (fm). When it passes through AOM, the frequency becomes fdiff = fm ±80 MHz. If this frequency beam and the PSL laser beam is aligned properly, and if this frequency is equal to the product of an integer and the free spectral range of the cavity, this will resonate in the cavity.  Then we expect to get a peak in the ETM transmission spectrum corresponding to the frequency we injected through the optical Fibre.

Through out the experiment we need to make sure that the PSL is locked. Thus, the signal detected by the photo detector when only PSL is resonating inside the cavity, act as a DC signal. Then we give a narrow scan to the Marconi. When fdiff = N*FSRy this condition is satisfied, we will observe a peak in the output. Here FSRy  is the free spectral range of the cavity which is approximately equal to 3.893 MHz.

Yesterday afternoon, Johannes, Koji and myself tried to observe this peak. We aligned the cavity by observing the output signal from the AS100 photo detector. We made the alignment in such a way that the intensity output getting from this photo detector is maximum. We used a Spectrum analyser to see the output. After that we connected a photo detector to collect the YEND transmission signal from the ETM mirror. We used a lens to focus this directly to the photodetector. Then we connected this photodetector to the spectrum analyser, which was located near the AS table. We took a large cable to meet this purpose. But still the cable was not lengthy enough, so we joined it with another cable and finally connected it with the spectrum analyser. Then we gave a scan to the Marconi from 51 MHZ to 55 MHz. We repeated this experiment with a scan of 55 MHz to 59 MHz also. We repeated this a few times, but we were not able to see the peak.

We assume that this can be because of some issue with the alignment or it can be because of some issue with the photo detector we used. We would like to repeat this experiment and get the signal properly.

I am attaching a flow chart of the setup and also a picture of the mirrors and photo detector we inserted in the Y-End table.

 
Attachment 1: photodetector_alignment.jpg  3.688 MB  Uploaded Thu Jun 7 11:55:29 2018  | Hide | Hide all
photodetector_alignment.jpg
Attachment 2: design1.PNG  21 kB  Uploaded Thu Jun 7 11:55:49 2018  | Hide | Hide all
design1.PNG
ELOG V3.1.3-